A small ball of mass O, 5 kg hangs on a non-stretchable cord. The cord is placed horizontally and then the ball is released. Find the tension force of the cord when it forms angle 30° with the vertical. The answer should be expressed in newtons [N]. Assume the free fall acceleration to be 10 m/s^2.


1
Expert's answer
2020-11-25T10:51:45-0500

Find the speed of the ball when the cord (length "l") makes a 30° angle:


"v=\sqrt{2gh}=\sqrt{2gl\text{ sin}30°}."

Find the centripetal acceleration at this speed:


"a=\frac{v^2}{l}=2g\text{ sin}30°."

The acceleration creates a force of


"F=ma=2mg\text{ sin}30°."

The force of gravity that acts on the cord at this angle is


"F_g=mg\text{ cos}60°."

So, the tension is


"T=F+F_g=\\\\=mg[\\text{cos}(90\u00b0-30\u00ad\u00b0)+2\\text{sin30\u00b0}]=7.5\\text{ N}."